OpenStudy (anonymous):

Is this correct? Nathan opened a savings account 11 years ago with a deposit of $2,665.79. The account has an interest rate of 4.8% compounded quarterly. How much interest has Nathan earned? $2,793.75 $1,798.99 $1,839.97 $4,505.76 <-- My answer (:

hero (hero):

Compound Quarterly: \[A = P\left(1 + \frac{r}{4}\right)^{4t}\]

OpenStudy (anonymous):

That's the formula I used

hero (hero):

Hang on a minute

OpenStudy (anonymous):

Okay, will do (:

hero (hero):

It is possible you may have mis-interpreted the question

OpenStudy (anonymous):

I tend to do that a lot, so it's not surprising if I did lol

OpenStudy (anonymous):

Can I show how I set it up to see if I did it wrong?

hero (hero):

Yes, to find the interest you have to do this Subtract A - P, so 4505.76 - 2665.79 = 1839.97

hero (hero):

They only want the interest gained, not the full value of A.

OpenStudy (anonymous):

Okay, I definitely misinterpreted it. I just don't really know how to find how much interest was earned

hero (hero):

I just showed it to you.

hero (hero):

I = A - P I = Interest Earned P = Principal A = Account Balance

OpenStudy (anonymous):

Oh, sorry. Okay, gimme a minute and I'll show you what I ended up with

OpenStudy (anonymous):

Well, I ended up with the same answer as you. Does that mean I was correct?

hero (hero):

Yes, 1839.97 is the correct answer for this.

OpenStudy (anonymous):

but it is not one of my answer choices :l

OpenStudy (anonymous):

This is what I have to choose from: Nathan opened a savings account 11 years ago with a deposit of $2,665.79. The account has an interest rate of 4.8% compounded quarterly. How much interest has Nathan earned? $2,793.75 $1,798.99 $1,839.97 $4,505.76

hero (hero):

That's interesting because it is one of the answer choices.

OpenStudy (anonymous):

...I can be so stupid and oblivious at times.........

hero (hero):

Nathan opened a savings account 11 years ago with a deposit of $2,665.79. The account has an interest rate of 4.8% compounded quarterly. How much interest has Nathan earned? $2,793.75 $1,798.99 \(\color\red{$1,839.97}\) $4,505.76

OpenStudy (anonymous):

Thanks Hero. As embarrassing as it was, thanks for your help

hero (hero):

You're welcome

OpenStudy (rockinhood):

Lol Savage @Hero